OpenStudy (anonymous):

How would I solve for x? 9^x-1=27

OpenStudy (anonymous):

\[9^{x-1} = 27\]

OpenStudy (luigi0210):

Get common bases: \[3^{2(x-1)}=3^3\] \[2x-2=3\] Finish it

OpenStudy (anonymous):

Oh okay thank you!

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

I would use that same method with 3^2x+2= 8 right? @Luigi0210

OpenStudy (luigi0210):

You might have to use logs, or just graph them.

OpenStudy (anonymous):

How would I use logs on this?

OpenStudy (luigi0210):

\[3^{2x}=6\] equals: \[\log_{3} 6=2x\]

OpenStudy (anonymous):

where did the 6 come from

OpenStudy (anonymous):

i thought it would be 4

OpenStudy (luigi0210):

on the 3^2x+2=8 right? just subtract 2

OpenStudy (anonymous):

\[3^{2x+2}=8\] i thought you would divide it but i guess i'm not sure

OpenStudy (luigi0210):

Oh the numbers are all exponents? You should of mentioned that in the beginning :P That as a log would look like this: \[\log_{3} 8=2x+2\]

OpenStudy (anonymous):

oh okay thanks and then how would i take the log of that? I'm new to logs so I'm really not sure what I'm doing

OpenStudy (luigi0210):

Then just subtract 2 and divide, make sure to not convert the log just yet because accuracy counts: \[\frac{\log_{3} 8-2}{2}=x\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

log38-1=x?

OpenStudy (anonymous):

\[\log _{3}8-1=x\]

OpenStudy (luigi0210):

No, you have to do it all in one calculation..

OpenStudy (anonymous):

how

OpenStudy (luigi0210):

You know, it might easier to graph this one for you ^_^' The points of intersection are where your answer is at

OpenStudy (anonymous):

thank you ! if i plugged in the equation in my calculator would it give me the same results?

OpenStudy (luigi0210):

It should